OpenStudy (anonymous):

What are the next three terms in the sequence 4096, 1024, 256, 64, ...? 16, 4, 0 16, 4, 1 32, 8, 2 32, 8, 4

OpenStudy (anonymous):

it is a G.P. with c.r=1024/4096=1/4 a,b,c,d are in G.P. b=ar,c=br,d=cr,....

OpenStudy (anonymous):

???

OpenStudy (anonymous):

16, 4, 1 Each number is divisible by 4 from the preceding number. xx

OpenStudy (anonymous):

ok thx